# Dark Matter Substructure or Source Model Systematics? A Case Study of Cluster Lens Abell S1063

### v1

@@ 5 Discussion & Conclusions
At the same time, CAB is a local formalism that should be used with caution. Overestimating the region of validity can lead to the failure of the CAB formalism, leading to lens modeling systematics that similarly bias the inferred parameters. A multi-band analysis in different mask sizes allows us to detect the failure of the local lensing formalism as a tension between lens parameters across filters. We demonstrate that CAB works better in smaller masks, as seen in reduced tension between parameters from different filters. Hence, we show the power of multi-band analysis as an empirical approach to determine the region of validity of this local formalism.
Source and lens modeling systematics have profound implications for substructure search. Our statistically significant detection of a subhalo with shapelet source modeling tells a cautionary tale that source modeling systematics can lead to convincing substructure detections. This is caused by insufficient complexity in the source model, leading to a preference for a subhalo that accounts for some of the missing complexity in the source-light distribution. At the same time, the higher level of lens modeling systematics seen in Large Masks exacerbates spurious detections. As with the source model, a subhalo can likewise account for some of the missing complexity in the lens model. Multi-band analysis is key to disentangling source and lens modeling systematics from substructure detection. While seemingly convincing, several qualities give away hints that the perturber detected with shapelets may be spurious.
@@
The statistical significance of the subhalo detection diminishes as we complicate the source model by increasing the shapelet order parameter .
By introducing pixel-based source modeling with Delaunay triangulation, we allow for a highly flexible source model without any explicit regularization. With this higher source complexity, the preference of our model for substructure disappears. Interestingly, when imposing regularization on Delaunay source modeling, spurious substructure appears again, although at different location, in two filters. This finding shows that strong regularization conditions can lead to spurious substructure detections, as additional complexity in the lens model accounts for the limited complexity in source modeling. This result is particularly relevant for JWST cluster lenses, where multiply-lensed, magnified images offer remarkably high resolutions, and a highly detailed source model becomes essential. However, we expect the result to translate to galaxy-galaxy strong lenses, where strong regularization conditions that limit the source model complexity could lead to a spurious preference for substructure that accounts for some of this missing complexity. The availability of multi-filter data with JWST gives us a higher chance of decisively ruling out spurious detections.
@@ Acknowledgements
We thank Arthur Tsang for helpful discussions. CD is partially supported by the Department of Energy (DOE) Grant No. DE-SC0020223.
@@ Appendix A Posteriors from Large Masks
We show the full posterior distribution of the Curved Arc Basis parameters obtained with shapelet and Delaunay source modeling methods, using Large Masks, in Figures 21 and 22.
@@ Appendix B Posteriors from Small Masks
We show the full posterior distribution of the CAB parameters obtained with shapelet and Delaunay source modeling methods, using Small Masks, in Figures 23 and 24. We compare the means and errors from the two source modeling methods in Figure 25.
@@ Appendix C Spurious subhalo statistics
Appendix D Mock Subhalo Tests
To compare the performance of shapelet and Delaunay source modeling methods in detecting substructures in real data, we test subhalo recovery in realistic mocks of JWST quality. We lens a real high-resolution source NGC 1300, a barred spiral galaxy, with smooth Curved Arc Basis deflections and a single subhalo in the line of sight of one of the lensed images. We choose a subhalo mass of , concentration of , position of (, ) = (", ") relative to the brightest lensed image pixel, at the same subhalo redshift of as the Abell S1063 cluster. The brightness, sizes, noise levels, and Point Spread Functions of the mocks are identical to our JWST observations of system 4 in filter F115W.
We fit the mock images with shapelet and Delaunay source modeling methods, using the same masks. In shapelet source modeling, we minimized the BIC of the residuals using the true CAB and NFW parameters to find the optimal shapelet order . In Delaunay source modeling, we pick half of the unmasked pixels in the first mock image as candidate tesselation points, as described in Section 3.2.2. If any of the chosen candidate points end up being two neighboring pixels, one of them is removed.
We show our mock images, model reconstructions, and residuals from Delaunay and shapelet source modeling in Figure 32, as well as the posterior distribution of the CAB and NFW parameters in Figure 33. We find that Delaunay is consistently less biased than shapelets in measuring the smooth Curved Arc Basis parameters, which contributes to a more accurate measurement of the subhalo mass and concentration. We conclude that Delaunay is a reliable source modeling method for a robust substructure search, especially in lensing systems involving source galaxies with complex morphologies.

### v2

@@ 5 Discussion & Conclusions
At the same time, CAB is a local formalism that should be used with caution. Overestimating the region of validity can lead to the failure of the CAB formalism, leading to lens modeling systematics that similarly bias the inferred parameters. A multi-band analysis in different mask sizes allows us to detect the failure of the local lensing formalism as a tension between lens parameters across filters. We demonstrate that CAB works better in smaller masks for all source reconstruction methods, as seen in reduced tension between parameters from different filters. Hence, we show the power of multi-band analysis as an empirical approach to determine the region of validity of this local formalism.
Source and lens modeling systematics have profound implications for substructure search. Our statistically significant detection of a subhalo with Cartesian shapelets tells a cautionary tale that source modeling systematics can lead to convincing substructure detections. Switching to elliptical shapelets weakens the statistical preference for the spurious subhalo. This suggests that insufficient complexity of the source model can lead to a preference for a subhalo that accounts for some of the missing complexity in the source-light distribution. At the same time, the higher level of lens modeling systematics seen in Large Masks exacerbates spurious detections. As with the source model, a subhalo can likewise account for some of the missing complexity in the lens model. Multi-band analysis is key to disentangling source and lens modeling systematics from substructure detection. While seemingly convincing, several qualities give away hints that the perturber detected with shapelets may be spurious.
@@
The statistical significance of the subhalo detection diminishes as we complicate the source model by increasing the shapelet order parameter or allowing for ellipticity of the shapelet basis.
By introducing pixel-based source modeling with Delaunay triangulation, we allow for a highly flexible source model without any explicit regularization. With this higher source complexity, the preference of our model for substructure disappears. Interestingly, when imposing gradient regularization on Delaunay source modeling, spurious substructure appears again, although at different location, in two filters. This finding demonstrates that strong, non-adaptive regularization conditions can lead to spurious substructure detections, as additional complexity in the lens model accounts for the limited complexity in source modeling. This result is particularly relevant for JWST cluster lenses, where multiply-lensed, magnified images offer remarkably high resolutions, and a highly detailed source model becomes essential. However, we expect the result to translate to galaxy-galaxy strong lenses, where strong regularization conditions that limit the source model complexity could lead to a spurious preference for substructure that accounts for some of this missing complexity. The availability of multi-filter data with JWST gives us a higher chance of decisively ruling out spurious detections.
@@ Acknowledgements
We thank Arthur Tsang for helpful discussions. CD is partially supported by the Department of Energy (DOE) Grant No. DE-SC0020223. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program .
@@ Appendix A Posteriors from Large Masks
We show the full posterior distribution of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling methods, using Large Masks, in Figures 21 and 22.
@@ Appendix B Posteriors from Small Masks
We show the full posterior distribution of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling methods, using Small Masks, in Figures 23 and 24. We compare the means and errors from the two source modeling methods in Figure 25.
@@ Appendix C Spurious subhalo statistics
Appendix D Elliptical Shapelets
We repeat the smooth CAB and substructure analyses using an elliptical instead of Cartesian shapelet basis. We show the resulting posterior distributions from CAB analyses in Large and Small Masks in Figures 32 and 33, respectively. We show the BIC statistic of smooth and substructure CAB models for a range of shapelet-order parameters in Figures 34, 35, and 36. We compare the statistical significance of subhalo detection between analyses relying on Cartesian and elliptical shapelet bases for source reconstruction in Figures 37 and 38.
Appendix E Mock Subhalo Tests
To compare the performance of shapelet and Delaunay source modeling methods in detecting substructures in real data, we test subhalo recovery in realistic mocks of JWST quality. We lens a real high-resolution source NGC 1300, a barred spiral galaxy, with smooth CAB deflections and a single subhalo in the line of sight of one of the lensed images. We choose a subhalo mass of , concentration of , position of (, ) = (", ") relative to the brightest lensed image pixel, at the same subhalo redshift of as the Abell S1063 cluster. The brightness, sizes, noise levels, and Point Spread Functions of the mocks are identical to our JWST observations of system 4 in filter F115W.
We fit the mock images with Cartesian shapelet and Delaunay source modeling methods, using the same masks. In shapelet source modeling, we minimized the BIC of the residuals using the true CAB and NFW parameters to find the optimal shapelet order . In Delaunay source modeling, we pick half of the unmasked pixels in the first mock image as candidate tesselation points, as described in Section 3.2.2. If any of the chosen candidate points end up being two neighboring pixels, one of them is removed.
We show our mock images, model reconstructions, and residuals from Delaunay and Cartesian shapelet source modeling in Figure 39, as well as the posterior distribution of the CAB and NFW parameters in Figure 40. We find that Delaunay is consistently less biased than Cartesian shapelets in measuring the smooth CAB parameters, which contributes to a more accurate measurement of the subhalo mass and concentration. We conclude that Delaunay is a reliable source modeling method for a robust substructure search, especially in lensing systems involving source galaxies with complex morphologies.